# Undergraduate Course Allocation

### v1

@@ Appendix B Proof of Theorem 1 (Omitted Details)
$z_{S\backslash\{1,...,L^{\prime}\}}(t^{*})+\sum_{s=1}^{L^{\prime}}\sum_{f=1}^{w_{s}}a_{s}^{f}d^{f}_{s}$
where $0\leq a_{s}^{f}\leq 1,s=1,...,L^{\prime},f=1,...,w_{s}$ and $\sum_{f=1}^{w_{s}}a_{s}^{f}=1,$ $s=1,...,L^{\prime}$. Budish (2011) shows in Step 8 of Theorem 1 that there exists an element of the above geometric structure, $z(t^{\phi^{\prime}})$, within $\sqrt{kM/2}$ distance from the perfect market-clearing excess demand vector. Step 9 explains how to adjust agent budgets to find an approximate competitive equilibrium at $t^{*}$ with excess demands equal to $z(t^{\phi^{\prime}})$. These purely mathematical arguments remain unchanged from the original paper.∎
@@ Appendix C The Student Utility Model Calibration
This appendix provides details on the method of simulated moments approach used to calibrate the student utility model. Section C.1 outlines the student utility model. Section C.2 discusses the necessary adjustments to the university’s course reserves. Section C.3 presents the empirical moments we use to calibrate the model. Section C.4 outlines the iterated process we use to calibrate the model and presents the calibration results. Section C.5 provides details on model fit.
@@ C.1 Model
  • $Y_{s}$ and $H_{c}$ are vectors of student $s$’s observed characteristics and course $c$’s observed characteristics, respectively. Both vectors are binary vectors having all but one element as zero. Denote the set of colleges by ${\cal A}=\{A,...,G\}$ and years of study by ${\cal Y}=\{1,...,4\}$. Then, $Y_{s}\in\{0,1\}^{|{\cal A}||{\cal Y}|}$ determines the college and year of study of student $s$ and $H_{c}\in\{0,1\}^{|{\cal A}|}$ determines the college of course $c$.
@@
We assume that the utility from taking multiple courses equals the sum of utilities of taking individual courses. The utility of taking no course is normalized to $u_{s0}=0$. Each student can be enrolled in up to a maximum of five courses.
The parametric form for student utilities in equation (C1) has three non-random components. The first, $Y^{\prime}_{s}\theta$ is a horizontal component that is determined solely by the student’s characteristics. The vector of coefficients $\theta\in\mathbb{R}^{|{\cal A}||{\cal Y}|}$ that has $28$ elements in total to calibrate. The horizontal component determines the importance of taking courses for student $s$, with a larger value meaning student $s$ wants to enroll in more courses. This accounts for variability in the data in the number of courses taken by students from different colleges and years of study, as shown in Table C.1.
The second component, $Y^{\prime}_{s}\Gamma H_{c}$, is an interaction term that determines the heterogeneity of students’ preferences over courses. We restrict the matrix $\Gamma$ to have $49$ independent parameters $\gamma_{aa^{\prime}}$, where $a$ is student $s$’s college and $a^{\prime}$ is course $c$’s college. The matrix $\Gamma$ accounts for heterogeneity in patterns of course enrollment across colleges observed in the data demonstrated in Table C.2. For example, students from college $A$ take $71\%$ of courses from their home college, while more than $95\%$ of courses taken by students from college $C$ are courses outside their home college.
The third component $z_{c}$ is a vertical component that determines the attractiveness of course $c$. This component is common among students and determines the quality of the course. The number of vertical components coincides with the number of courses, $756$.
@@
Notes: Each row of this table presents the average number of courses taken by students from the same college and different years of study.
Table C.2: The percentages of courses taken by students across colleges
@@
A 71 0.47 1.9 17 1.7 5 2.7
B 1.9 44 0.61 18 19 12 4.1
C 10 1.3 4.5 37 29 15 2.4
D 2.5 0.98 0.83 64 12 13 6.8
E 1.2 2.6 0.14 24 54 16 2.1
F 1.5 1.8 0.84 23 18 52 2.3
G 1.3 0.39 0.26 31 7.7 6.1 53
Notes: The table the percentage of courses taken by students across colleges. The rows correspond to students’ colleges, and the columns correspond to courses’ colleges. For example, the cell in row $2$ and column $1$ shows that $1.9\%$ of courses taken by students from college $B$ are courses from college $A$.
Parameter normalization. To correctly interpret the interaction term $\Gamma$ and the vertical component $z$, we need to further normalize the model’s parameters. As described above, for student $s$ from college $a$ and course $c$ from college $a^{\prime}$, we have $Y^{\prime}_{s}\Gamma H_{c}=\gamma_{aa^{\prime}}$. Let $S_{a}$ and $S_{ay}$ be the set of students from college $a$ and the set of students from college $a$ and year $y$, respectively. The average utility of students from college $a$ from taking course $c$ from college $a^{\prime}$ equals
@@
One possible normalization is $\overline{\theta}_{a}=0$ and $\gamma_{aa}=0$ for $a=1,...,7$. This normalization leads to the interpretation of $z_{c}$ as the average popularity of the course among students from the same college. Then, $\gamma_{aa^{\prime}}$ is an increase in student’s utility from taking a course in college $a^{\prime}$ compared to a course in student’s college $a$ with the same vertical component.
Students’ choice sets. The student choice sets are unobserved in the data. We do not observe students’ past enrollments, and, as a result, we do not know each student’s set of courses for which she satisfies course prerequisites. In addition, students’ choice sets are influenced by the courses chosen by students with earlier time-slots in the course allocation process, and, hence, students’ choice sets are determined endogenously. There are further complications associated with the changing of course reserves during the course allocation process, which are described in Section C.2.
These complications require us to take a stance on how student’s choice sets are formed. We assume that each student’s choice set is limited to $80$ courses. The courses in the student’s choice set are drawn randomly such that the probability that a course from college $a^{\prime}$ is drawn equals the share of students from the same college-year who are enrolled in courses in college $a^{\prime}$ in the actual data. Additionally, we make adjustments for courses with course reservations. For each course reservation with $r$ reserved seats, we ensure that the course is included in the choice sets of at least $2r$ students who qualify for the reservation. This adjustment prevents course reservations from artificially limiting a course?s capacity due to an insufficient number of qualified students. [^31] If there are fewer than $2r$ qualified students, the course is included in the choice set of all such students.
Random choice sets are a limitation of our model. Any changes in the way we draw the course lists could potentially change the calibrated parameter values. Crawford, Griffith and Iaria (2021) provide an insightful discussion of preference estimation techniques with unobserved choice sets. Still, the size of the student choice sets we consider is reasonable, as previous studies report that students typically consider only a limited subset of courses in a given semester (see Budish and Cantillon, 2012; Diebold and Bichler, 2017). We explore how the performance of the Pseudo-Market with Priorities mechanism changes for various sizes of choice sets (and the size of the noise parameter) using simulations in Appendix D.4.
@@ C.2 Course Reserve Adjustment
We use the method of simulated moments to calibrate the student utility model parameters. We need, however, to perform some adjustments to course reserves. Course reserves are typically set large at the beginning of the course allocation process and relaxed at later stages. Hence, the final course allocation can violate the initial course reserves. In fact, $170$ out of $756$ courses violate initial course reservations in the data.
We next describe the necessary and sufficient conditions for the existence of a matching that assigns a given population of students to the course and satisfies the course reservation criteria. Afterward, we describe the process of adjusting course reserves to ensure that the number of reserved seats is consistent with the course seat assignment observed in the data.
Two course reserve criteria. Consider a course with a capacity of $q$ seats. Denote the number of course reserves for students who satisfy criterion $A_{i}$ (student’s department and year of study) as $\mathfrak{r}_{A_{i}}$, $i=1,2$. We assume that $\mathfrak{r}_{A_{1}}\geq 0$ or $\mathfrak{r}_{A_{2}}\geq 0$ and $\mathfrak{r}_{A_{1}}+\mathfrak{r}_{A_{2}}\leq q$. One should interpret the reservation such that if a student satisfies both $A_{1}$ and $A_{2}$, she is eligible to take a seat satisfying either criterion, but if she takes $A_{1}$-type seats it does not decrease the number of $A_{2}$-type seats. This interpretation is consistent with the actual course enrollment process in the data. The number of students who satisfy criterion $A_{i}$ and who are matched to the course in the final allocation is denoted by $x_{A_{i}}$, $i=1,2$. Students who do not satisfy any course reserve criteria are called 00-students, and we denote the number of such students assigned to the course as $x_{0}$. Then, the course assignment satisfies course reserves if and only if
@@
To state the theorem, let us consider a set of students $S$ and course reserves criteria $A_{1},...,A_{I}$ for some $I\geq 1$. The number of reserved seats corresponding to criterion $A_{i}$ equals $\mathfrak{r}_{A_{i}}\geq 0$. We denote the set of available course seats as $Y$ with $|Y|=q$ and assume $\sum_{i=1}^{I}\mathfrak{r}_{A_{i}}\leq q$. Each student could potentially satisfy multiple criteria. We consider a bipartite graph $(S,Y,E)$ where $E$ is the set of edges connecting student and seats. A student $s\in{\cal S}$ is connected to seat $y\in Y$ with an edge if student $s$ is eligible to take seat $y$. For example, if there are no course reserves, each student is connected with an edge to each seat ($q$ edges). If there is one course reservation criterion $A$, each 00-student has $q-\mathfrak{r}_{A}$ edges, and each student who satisfies criterion $A$ has $q$ edges.
For a subset of students $S^{\prime}\subseteq S$, let $N(S^{\prime})$ denote the set of seats in $Y$ that are connected to at least one student in $S$. We call these seats the neighbors of $S^{\prime}$. A perfect matching is a matching between students and seats in which each student is assigned exactly one seat. We are now ready to state the theorem.
@@
  • The number of students who satisfy no criteria, that is, 00-students
@@
  • The number of students who satisfy only criterion $A_{i}$ or no criteria (adding 00-students only increases the number on the left-hand side of inequality (C3) and does not change the number on the right-hand side of the inequality)
@@
  • Let us now denote =${\cal I}=\{1,...,I\}$ and take some subset of available course reserve criteria ${\cal I^{\prime}}\subseteq{\cal I}$. We consider the number of students who satisfy at least one criterion in the set $\{A_{i}:i\in{\cal I^{\prime}}\}$ or no criteria and who do not satisfy the criteria in the complementary set $\superscript\{A_{i}:i\in{\cal I}\backslash{\cal I^{\prime}}\}$. Hall’s Marriage Theorem implies that the number of these students must satisfy
@@
Overall, Hall’s Marriage Theorem implies that the set of inequalities (C4) for all possible subsets of criteria ${\cal I^{\prime}}\subseteq{\cal I}$ provides the necessary and sufficient conditions for the existence of a matching of students to course seats that satisfies the course reserve criteria.
@@ C.3 Moment Conditions
For each set of parameter values and the realization of random components of the student utility model described in Section C.1, we use the Random Serial Dictatorship with adjusted course reserves to obtain the allocations of courses to students. Then, we use these allocations to calculate the simulated moments specified below.
@@
Courses at maximum capacity. A critical aspect of the data is that $85$ out of $756$ courses are at maximum capacity in the data, which creates a problem with identifying course-specific parameters $z_{c}$ for such courses. To address this issue, we replace moments $\tilde{m}_{c}$ for such courses with other moments based on the timing of course enrollment. We discuss these moments in more detail below.
To address the problem of identifying the vertical components $z_{c}$ for courses $c\in{\cal C}$ at maximum capacity, we use a crucial aspect of the data—the timing of student course enrollment. To track student course enrollment, we first define student ranking.
@@
Hence, more popular courses are filled up earlier with students of a smaller rank. Less popular courses are filled up later with students of a larger rank. [^33] More senior students have earlier time slots in the registration process. Hence, the measuring popularity by how fast a course fills up is biased towards the popularity of courses among more senior students.
@@
For courses at maximum capacity, we replace moments $\tilde{m}_{c}$ with the mean rank across the registered students. [^35] We also tried to use the largest student rank and some quantiles of students registered for the course to measure course popularity. These measures did not work very well. Last-minute course registrations are a noisy measure that do not reflect course popularity well. Measures based on quantiles proved to be discontinuous, with a small change in utility parameters often resulting in jumps in the student ranks at a specific quantile and a bad model fit. We keep using the symbol $\tilde{m}_{c}$ for all moments corresponding to vertical components $z$ to avoid clutter in notation.
Overall, there are $756+28+49=833$ moments $m=(\tilde{m},\breve{m},\hat{m})$ and $819$ independent parameters $\pi=(\theta,\Gamma,z)$ to calibrate. Hence, the model is over-identified, allowing us to test for model fit in Section C.5.

### v2

@@ Appendix B Proof of Theorem 1 (Omitted Details)
[…]
where $0\leq a_{s}^{f}\leq 1,s=1,...,L^{\prime},f=1,...,w_{s}$ and $\sum_{f=1}^{w_{s}}a_{s}^{f}=1,$ $s=1,...,L^{\prime}$. Budish (2011) shows in Step 8 of Theorem 1 that there exists an element of the above geometric structure, $z(t^{\phi^{\prime}})$, within $\sqrt{kM/2}$ distance from the perfect market-clearing excess demand vector. This purely mathematical argument remains unchanged.
Step 9. While $z(t^{\phi^{\prime}})$ is within $\sqrt{kM/2}$ distance from the perfect market-clearing excess demand vector, it may be that $t^{\phi^{\prime}}\in\tilde{\mathcal{T}}\setminus\mathcal{T}$ rather than $\mathcal{T}$. Since we know that $t^{*}\in\mathcal{T}$, we can alter student budgets $b$ so that excess demand at $t^{*}$ equals $z(t^{\phi^{\prime}})$.
For any student $s\in\mathcal{S}\setminus\{1,...,L^{\prime}\}$, we know that $s$ selects the same course schedule at $t^{\phi^{\prime}}$ and $t^{*}$. As a result, setting $b^{*}_{s}=b_{s}+\tau^{\prime}_{s,x^{*}_{s}}$, the price of any bundle preferred by $s$ to $x^{*}_{s}$ costs more than $b^{*}_{s}$.
Any student $s\in\{1,...,L^{\prime}\}$ may select a different course schedules at $t^{\phi^{\prime}}$ and $t^{*}$. For each such student, for each intersecting budget constraint $m=1,...,w_{s}$, if $t^{\phi^{\prime}}\in H^{1}(s,x^{m}_{s})$, then the course schedule $x^{m}_{s}$ is unaffordable for student $s$ at $t^{\phi^{\prime}}$. To guarantee that $x^{m}_{s}$ is unaffordable at $t^{*}$ as well, as in Budish (2011), we change the reverse tax $\tau^{\prime}_{s,x^{m}_{s}}$ to be $\tau^{\prime\prime}_{s,x^{m}_{s}}=\tau^{\prime}_{s,x^{m}_{s}}-\delta_{2}$, where $\delta_{2}>0$ is small enough to ensure that conditions (i), (ii), and (iii) from Lemma B1 still hold. If $t^{\phi^{\prime}}\in H^{0}(s,x^{m}_{s})$ (or, for any bundle $x\neq x^{m}_{s}$ for $m=1,...,w_{s}$), set $\tau^{\prime\prime}_{s,x^{m}_{s}}=\tau^{\prime}_{s,x^{m}_{s}}$. Taxes still favor more preferred bundles, i.e., $\tau^{\prime\prime}_{s,x_{s}}>\tau^{\prime\prime}_{sx^{\prime}_{s}}$ if $x^{\prime}_{s}\succ_{s}x_{s}$. Then, letting $x^{*}_{s}=d_{s}(t^{*},b_{s},\tau^{\prime\prime}_{s})$ and setting $b^{*}_{s}=b_{s}+\tau^{\prime\prime}_{s,x^{*}_{s}}$, the price of any bundle preferred by $s$ to $x^{*}_{s}$ costs more than $b^{*}_{s}$.
As a result, with all taxes equal to 0, at budgets $b^{*}$ and priority-specific course prices $p^{*}$ derived from $t^{*}$ using equation (A1), the resulting allocation $x^{*}$ generates excess demand equal to $z(t^{\phi^{\prime}})$, which has a market-clearing error no greater than $\sqrt{kM/2}$. ∎
@@ Appendix C The Student Utility Model Calibration
This appendix describes the method of the simulated moments approach used to calibrate the student utility model. First, Section C.1 outlines the student utility model. Then, Section C.2 discusses adjustments to the university’s course reserves data. The empirical moments used to calibrate the model are described in Section C.3. Sections C.4 and C.5 provide the calibration process, calibration results, and details on model fit.
@@ C.1 Model
  • $Y_{s}$ and $H_{c}$ are binary vectors of student $s$’s observed characteristics and course $c$’s observed characteristics, each with one nonzero element. Denote the set of colleges by ${\cal A}=\{A,...,G\}$ and years of study by ${\cal Y}=\{1,...,4\}$. Then, $Y_{s}\in\{0,1\}^{|{\cal A}||{\cal Y}|}$ sets the college and year of study of student $s$ and $H_{c}\in\{0,1\}^{|{\cal A}|}$ sets the college of course $c$.
@@
Each student can be enrolled in up to a maximum of five courses. We assume that the utility from taking multiple courses equals the sum of utilities of taking individual courses. The utility of taking no course is normalized to $u_{s0}=0$.
The parametric form for student utilities in equation (C1) has three non-random components. The first, $Y^{\prime}_{s}\theta$ is a horizontal component that is determined solely by the student’s characteristics. The vector of coefficients $\theta\in\mathbb{R}^{|{\cal A}||{\cal Y}|}$ has $28$ elements in total to calibrate. The horizontal component determines the importance of taking courses for student $s$, with a larger value meaning student $s$ wants to enroll in more courses. This accounts for variability in the data in the number of courses taken by students from different colleges and years of study, as shown in Table C.1.
The second component, $Y^{\prime}_{s}\Gamma H_{c}$, is an interaction term that determines the heterogeneity of students’ preferences over courses. We restrict the matrix $\Gamma$ to have $49$ independent parameters $\gamma_{aa^{\prime}}$, where $a$ is student $s$’s college and $a^{\prime}$ is course $c$’s college. $\Gamma$ accounts for the heterogeneity in patterns of course enrollment across colleges. For example, Table C.2 shows students from college $A$ take $71\%$ of courses from their home college, while more than $95\%$ of courses taken by students from college $C$ are courses outside their home college.
The third component $z_{c}$ is a vertical component that determines the attractiveness of course $c$. This component is common among all students.
@@
Notes: Each row presents the average number of courses taken by students from the same college and different years of study.
Table C.2: Percentages of courses taken by students across colleges
@@
A 71.4 0.5 1.9 16.9 1.7 5.0 2.7
B 1.9 43.8 0.6 18.4 19.0 12.2 4.1
C 10.2 1.3 4.5 37.2 29.2 15.2 2.4
D 2.5 1.0 0.8 63.9 12.0 13.0 6.8
E 1.2 2.6 0.1 24.1 53.7 16.1 2.1
F 1.5 1.8 0.8 23.3 17.9 52.4 2.3
G 1.3 0.4 0.3 31.3 7.7 6.1 53.0
Notes: This table reports the percentage of courses taken by students across colleges. The rows correspond to students’ colleges, and the columns correspond to courses’ colleges.
Parameter normalization. To correctly interpret the interaction term $\Gamma$ and the vertical component $z$, we need to further normalize the model’s parameters. As described above, for student $s$ from college $a$ and course $c$ from college $a^{\prime}$, we have $Y^{\prime}_{s}\Gamma H_{c}=\gamma_{aa^{\prime}}$. Let $S_{a}$ and $S_{ay}$ be the set of students in college $a$ and the set of students in college $a$ and year $y$, respectively. The average utility of students in college $a$ from taking course $c$ in college $a^{\prime}$ equals
@@
One possible normalization is $\overline{\theta}_{a}=0$ and $\gamma_{aa}=0$ for a=1,,7a=1,...,7. This normalization leads to the interpretation of $z_{c}$ as the average popularity of the course among students from the same college. Then, $\gamma_{aa^{\prime}}$ is an increase in student’s utility from taking a course in college $a^{\prime}$ compared to a course in student’s college $a$ with the same vertical component.
Students’ choice sets. The student choice sets are unobserved in the data. We do not track students’ past enrollments, and as a result, we do not know each student’s set of courses for which they satisfy course prerequisites. This complication requires us to take a stance on how students’ choice sets are formed. We assume that each student’s choice set is limited to $80$ courses. The courses in the student’s choice set are drawn randomly such that the probability that a course from college $a^{\prime}$ is drawn equals the share of students from the same college-year who are enrolled in courses in college $a^{\prime}$ in the actual data. Additionally, we make adjustments for courses with course reservations. For each course reservation with $r$ reserved seats, we ensure that the course is included in the choice sets of at least $2r$ students who qualify for the reservation. This adjustment prevents course reservations from artificially limiting a course’s capacity due to an insufficient number of qualified students. [^28] If there are fewer than $2r$ qualified students, the course is included in the choice set of all such students.
Random choice sets are a limitation of our model. Any changes in the way we draw the course lists could potentially change the calibrated parameter values. Crawford, Griffith and Iaria (2021) provide an insightful discussion of preference estimation techniques with unobserved choice sets. Still, the size of the student choice sets we consider is reasonable, as previous studies report that students typically consider only a limited subset of courses in a given semester (see Budish and Cantillon, 2012; Diebold and Bichler, 2017). We explore how the performance of mechanisms change for various sizes of choice sets in Appendix D.4.
@@ C.2 Course Reserve Adjustment
We use the method of simulated moments to calibrate the student utility model parameters. Some adjustments to the course reserves need to be performed first. Course reserves are typically set large at the beginning of the actual course allocation process and relaxed at later stages. Hence, the final course allocation can violate the initial course reserves. In fact, $170$ out of $756$ courses violate initial course reservations in the data.
We next describe the necessary and sufficient conditions for the existence of a matching that assigns a given population of students to a course and satisfies several reservation criteria. We begin by outlining the conditions for two reservation criteria and then extend these conditions to the general case. Afterward, we describe the process of adjusting course reserves to ensure that the number of reserved seats is consistent with the course seat assignment observed in the data.
Two course reserve criteria. Consider a course with a capacity of $q$ seats. Denote the number of course reserves for students who satisfy criterion $A_{i}$ (student’s department and year of study) as $\mathfrak{r}_{A_{i}}$, $i=1,2$. We assume that $\mathfrak{r}_{A_{1}}\geq 0$ or $\mathfrak{r}_{A_{2}}\geq 0$ and $\mathfrak{r}_{A_{1}}+\mathfrak{r}_{A_{2}}\leq q$. One should interpret the reservation such that if a student satisfies both $A_{1}$ and $A_{2}$, she is eligible to take a seat satisfying either criterion, but if she takes $A_{1}$-type seats it does not decrease the number of $A_{2}$-type seats. This interpretation is consistent with the actual course enrollment process in the data. The number of students who satisfy criterion $A_{i}$ and who are matched to the course in the final allocation is denoted by $x_{A_{i}}$, $i=1,2$. Students who do not satisfy any course reserve criteria are called 0-students, and we denote the number of such students assigned to the course as $x_{0}$. Then, the assignment satisfies course reserves if and only if
@@
To state the theorem, let us consider a set of students $S$ and course reserves criteria $A_{1},...,A_{I}$ for some $I\geq 1$. The number of reserved seats corresponding to criterion $A_{i}$ equals $\mathfrak{r}_{A_{i}}\geq 0$. We denote the set of available course seats as $Y$ with $|Y|=q$ and assume $\sum_{i=1}^{I}\mathfrak{r}_{A_{i}}\leq q$. Each student could potentially satisfy multiple criteria. We consider a bipartite graph $(S,Y,E)$ where $E$ is the set of edges connecting student and seats. A student $s\in{\cal S}$ is connected to seat $y\in Y$ with an edge if student $s$ is eligible to take seat $y$. For example, if there are no course reserves, each student is connected with an edge to each seat ($q$ edges). If there is one course reservation criterion $A$, each 0-student has $q-\mathfrak{r}_{A}$ edges, and each student who satisfies criterion $A$ has $q$ edges.
For a subset of students $S^{\prime}\subseteq S$, let $N(S^{\prime})$ denote the set of seats in $Y$ that are connected to at least one student in $S$. We call these seats the neighbors of $S^{\prime}$. A perfect matching is a matching between students and seats in which each student is assigned exactly one seat.
@@
  • The number of students who satisfy no criteria, that is, 0-students
@@
  • The number of students who satisfy only criterion $A_{i}$ or no criteria (adding 0-students only increases the number on the left-hand side of inequality (C3) and does not change the number on the right-hand side of the inequality)
@@
[…]
@@
[…]
@@ C.3 Moment Conditions
For each set of parameter values and realization of random components of the student utility model described in Section C.1, we use the Random Serial Dictatorship with adjusted course reserves to simulate the allocations of courses to students. Then, we use these allocations to calculate the simulated moments specified below.
@@
Courses at maximum capacity. A critical aspect of the data is that $85$ out of $756$ courses are at maximum capacity in the data, which creates a problem with identifying course-specific parameters $z_{c}$ for such courses. To address this issue, we replace moments $\tilde{m}_{c}$ for such courses with moments based on the timing of course enrollment. We first define a ranking of students.
@@
Hence, more popular courses are filled up earlier with students of a smaller rank. Less popular courses are filled up later with students of a larger rank. [^30] More senior students have earlier time slots in the registration process. Hence, measuring popularity by how fast a course fills up is biased towards the popularity of courses among more senior students.
@@
For courses at maximum capacity, we replace moments $\tilde{m}_{c}$ with the mean rank across the registered students. [^32] We also tried to use the largest student rank and some quantiles of students registered for the course to measure course popularity. These measures did not work very well. Last-minute course registrations are a noisy measure that do not reflect course popularity well. Measures based on quantiles proved to be discontinuous, with a small change in utility parameters often resulting in jumps in the student ranks at a specific quantile and a bad model fit. We keep using the symbol $\tilde{m}_{c}$ for all moments corresponding to vertical components $z$ to avoid clutter in notation. Overall, there are $756+28+49=833$ moments m=(m~,m˘,m^)m=($\tilde$$\breve{m}$$\hat${m}) and $819$ independent parameters $\pi=(\theta,\Gamma,z)$ to calibrate. Hence, the model is over-identified, allowing us to test for model fit in Section C.5.